OpenStudy (anonymous):

Write a two-step equation that needs the steps of subtraction then division in order to be solved, and then solve it. Show your work.

OpenStudy (anonymous):

2x+1=7 2x+1-1=7-1 2x=6 2x/2=6/2 x=3